OpenStudy (anonymous):

Suppose that A, B, C, are independent random variables, each being uniformly distributed over (0, 1). What is the probability that all of the roots of the equation Ax2 + Bx + C = 0 are real? P(B^2>= 4AC) = ?

OpenStudy (ryan123345):

Hey there! I see that you are new to OpenStudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! If you have any additional questions involving the website feel free to ask the people with the purples A's next to their name. They are the ambassadors and will answer all your questions regarding the site. And also read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Have a great day and I hope you have fun!

OpenStudy (kropot72):

E(A) = 0.5 E(C) = 0.5 \[E(A)\times E(C)=0.5\times 0.5=0.25\] \[4\times [E(A)\times E(C)]=4\times 0.25=1\] E(B) = 0.5 \[[E(B)]^{2}=0.5\times 0.5=0.25\] \[P(B ^{2}\ge 4AC)=\frac{0.25}{1}=\frac{1}{4}\]